An insurance company would like to test the hypothesis that a difference exists in the proportion of students in12th grade who text while driving when compared to the proportion of 11th

grade drivers who text. A random sample of 160 12th grade students found that 84 texted while driving. A random sample of 175 11th grade students found that 70 texted while driving. Assume Population 1 is defined as 12th-grade drivers and Population 2 is defined as 11th

grade drivers. What is the test statistic for this hypothesis test?

A.

2.29

B.

−1.73

C.

−2.47

D.

1.06
Answer #1

option A is correct
Answer #2

Answer:

Test Statistic: A.2.29


Steps:

  1. Sample Proportions:

    • p^1=84160=0.525 (12th grade).

    • p^2=70175=0.400 (11th grade).

  2. Pooled Proportion (p^):

    p^=84+70160+175=1543350.460

  3. Test Statistic (z):

    z=p^1p^2p^(1p^)(1n1+1n2)=0.5250.4000.460×0.540(1160+1175)2.29

Conclusion: The test statistic matches option A (2.29).


answered by: anonymous